OpenStudy (anonymous):

Yuri makes his own citrus vinaigrette salad dressing. He likes the dressing best if it contains 20% fruit juice. He has a bottle of dressing mixture that contains 50% juice and 3 cups of dressing that contains 10% juice. How much of the dressing mix with 50% juice should Yuri add to the dressing mix with 10% juice to get the dressing he prefers with 20% juice?

OpenStudy (anonymous):

So it is looking for a fraction of how much 50% mix Yuri should add to the 10% juice?

OpenStudy (anonymous):

That seems a bit strange.

OpenStudy (anonymous):

what does

OpenStudy (anonymous):

Adding 50% mix to the 10%. Because it states there are 3 cups of 10% juice.

OpenStudy (anonymous):

i know... im bad enough with word problems and this one i cant seem to understand.....

OpenStudy (calculusxy):

0.3 + 0.5x ----------------- = 0.2 3 + x Cross-multiply: 0.2(3 + x) = 0.3 + 0.5x 0.6 + 0.2x = 0.3 + 0.5x 0.3 = 0.3x x = 1 So...Yuri should add 1 cup of the 50% juice dressing to the 3 cups of the 10% juice dressing. He will then have 4 cups of 20% juice dressing. Source: Yahoo Answers

OpenStudy (calculusxy):

Did this help you?

OpenStudy (anonymous):

Yahoo answers comes to the rescue, xD

OpenStudy (anonymous):

lol...... yup thanks (: